the question is : you are given 1 point on a line and the slope of the line, find the coordinates of 3 other points on the line
(-2,-4), m = -2

Answer by ReadingBoosters(3246)

Apply the point-slope formula
y - y1 = m(x-x1)
...
y - -4 = -2(x - -2)
y + 4 = -2(x+2) ==> distribute the -2
y + 4 = -2x - 4 ==> isolate the y variable
y = -2x - 4 - 4
y = -2x - 8
...
choose any three values for x and solve for y, each one is a coordinate point
y = -2(1) - 8 = -2 - 8 = -10
(1, -10)
y = -2(3) - 8 = -6 - 8 = -14
(3, -14)
y = -2(-1) - 8 = 2 - 8 = -6
(-1, -6)
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.